OpenStudy (anonymous):

Brain Fart >.< 8 − x = 17

OpenStudy (anonymous):

Move over x to the other side So it will be like 8-x=17 -> 8-x+x=17+x so 8=x+17 should be easy from there

OpenStudy (anonymous):

where did you get 8-x+x=17+x ~confused

OpenStudy (anonymous):

I just added x on both sides :))

OpenStudy (anonymous):

so i can get rid of the -x;; if yu dont understand, just think of it as moving -x to the other side.

OpenStudy (anonymous):

i tried that when i tried answering the question my self when i added the x to the 17 ? >.<

OpenStudy (anonymous):

i dont care about the answwer i just want to know how you got that ahhh!!!

OpenStudy (anonymous):

8 - x = 17 (add x from both sides) 8 - x + x = 17 + x 8 = 17 + x (subtract 17 from both sides) 8 - 17 = 17 - 17 + x -9 = x

OpenStudy (anonymous):

^ there you go lol

OpenStudy (anonymous):

it can also be done like this... 8 - x = 17 (subtract 8 from both sides) 8 - 8 - x = 17 - 8 -x = 9 (multiply by -1 to make x positive) x = -9

OpenStudy (anonymous):

in the first one how did you get 8 - x + x = 17 + x

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

she add x to both sides

OpenStudy (anonymous):

Simply added (+x) on both sides. you can add a same number on both sides in math and it wont make any difference

OpenStudy (anonymous):

some people skip the writing part of this step, we are just showing you all the steps

OpenStudy (anonymous):

just remember when you are moving a number across the equal sign, the sign changes